OpenStudy (anonymous):

Divide then simplify by taking the roots. Assume that allexpressions under radicals represent positive numbers sqrt 324xy / 3 sqrt 3

OpenStudy (anonymous):

bet you didn't know that \[\sqrt{324}=18\] but it is true

OpenStudy (anonymous):

so you get \[\frac{18\sqrt{xy}}{3\sqrt{3}}\] \[=\frac{6\sqrt{xy}}{\sqrt{3}}\] \[=\frac{6\sqrt{xy}}{\sqrt{3}}\times \frac{\sqrt{3}}{\sqrt{3}}]=\frac{6\sqrt{3xy}}{3}=2\sqrt{3xy}\]

OpenStudy (anonymous):

thanks